# references.bib RefAV: Towards Planning-Centric Scenario Mining

### v2

@@ Abstract
Autonomous Vehicles (AVs) collect and pseudo-label terabytes of multi-modal data localized to HD maps during normal fleet testing. However, identifying interesting and safety-critical scenarios from uncurated driving logs remains a significant challenge. Traditional scenario mining techniques are error-prone and prohibitively time-consuming, often relying on hand-crafted structured queries. In this work, we revisit spatio-temporal scenario mining through the lens of recent vision-language models (VLMs) to detect whether a described scenario occurs in a driving log and, if so, precisely localize it in both time and space. To address this problem, we introduce RefAV, a large-scale dataset of diverse natural language queries that describe complex multi-agent interactions relevant to motion planning derived from driving logs in the Argoverse 2 Sensor dataset. We evaluate several referential multi-object trackers and present an empirical analysis of our baselines. Notably, we find that naively repurposing off-the-shelf VLMs yields poor performance, suggesting that scenario mining presents unique challenges. Lastly, we discuss our recent CVPR 2025 competition and share insights from the community. Our code and dataset are available on Github and Argoverse.
@@ 1 Introduction
Autonomous Vehicle (AV) deployment on public roads has increased significantly in recent years, with Waymo completing more than a million rides per week as of February 2025 [Korosec2025]. Despite the maturity of such autonomous ride-hailing services, AVs still get into accidents [thadani2023cruise] and often require manual intervention [korosec2025waymo, mihalascu2023cruise]. Although data-driven simulators are an integral component in establishing a safety case [karnchanachari2024towards, gulino2023waymax, vasudevan2024planning, yang2023unisim], validating end-to-end autonomy with real-world operational data remains a critical part of testing. However, identifying interesting and safety-critical scenarios from uncurated driving logs is akin to finding a "needle in a haystack" due to the scale of data collected during normal fleet operations. In this paper, we revisit the task of spatio-temporal scenario mining with recent vision-language models (VLMs) to identify interesting multi-agent interactions using natural language (cf. Fig. 1).
Status Quo. Although language-based 3D scene understanding has been extensively studied in the context of referential multi-object tracking (RMOT) [wu2023referkitti, wu2023nuprompt, li2025nugrounding], multi-modal visual question answering (VQA) [qian2024nuscenesqa, Inoue_2024_WACV, tian2025nuscenes], and VLM-based motion planning [sima2024drivelm, mao2023gpt, shao2024lmdrive, wang2024omnidrive], we argue that scenario mining presents unique challenges. RMOT extends referential grounding by associating referred objects over time. However, unlike RMOT, scenario mining does not guarantee that referred objects exist in a given driving log. Next, multi-modal VQA extends VQA with additional visual modalities like LiDAR. Although spatio-temporal scenario mining also leverages multi-modal input (e.g. RGB, LiDAR and HD maps), we require that methods output 3D tracks instead of text-based answers. Lastly, VLM-based motion planners directly estimate future ego-vehicle waypoints based on high-level language instructions and past sensor measurements. In contrast, scenario mining methods can reason over the full driving log to identify interactions between non-ego vehicles. Concretely, spatio-temporal scenario mining requires identifying if a described scenario occurs in a driving log from raw sensor measurements (e.g. RGB and LiDAR), and if so, precisely localizing all referred objects in both time and space with 3D tracks. To support this task, we propose RefAV, a large-scale dataset of diverse natural language queries designed to evaluate a model’s ability to find a visual “needle in a haystack”.
@@
Referential Tracking by Program Synthesis. Although prior referential trackers and VLMs achieve reasonable performance with simple referential prompts (e.g. find all cars), we find that such methods struggle with compositional reasoning and motion understanding (e.g. find all cars accelerating while changing lanes). To address this problem, we propose Referential Tracking by Program Synthesis (RefProg), a modular approach that combines off-the-shelf 3D tracks with LLMs. Inspired by recent work in program synthesis [surismenon2023vipergpt, gupta2023visual, marsili2025visual], our method uses an LLM to break down complex referential expressions into simpler compositional actions. Specifically, we define an API to describe hand-crafted atomic functions (e.g., turning, accelerating, changing lanes) and use an LLM to synthesize a program that composes these atomic actions corresponding to the natural language prompt. We then execute the generated program to filter off-the-shelf tracks to identify the subset of tracks that best align with the described scenario (cf. Fig. 3).
@@ 2 Related Works
Visual Question Answering with Vision-Language Models has improved significantly in recent years due to large-scale multi-modal pre-training. Vision-language models (VLMs) like LLaVA [liu2023llava, liu2023improvedllava], BLIP-2 [blip2], and Qwen2.5-VL [bai2025qwen2.5vl] show strong generalization across diverse domains [balanced_vqa_v2, tian2025nuscenes, qian2024nuscenesqa]. However, state-of-the-art models still struggle with spatial reasoning and grounding [madan2024revisiting, robicheaux2025roboflow100]. To address this issue, SpatialRGPT [cheng2024spatialrgpt] and SpatialVLM [chen2024spatialvlm] generate large-scale monocular depth pseudo-labels to improve 3D reasoning. Despite the effectiveness of such methods, zero-shot prompting yields poor compositional reasoning performance. Instead, VisProg [gupta2023visual] and ViperGPT [surismenon2023vipergpt] use LLMs generate executable code and use tools like OwlViT [minderer2022simple], CLIP [clip], MiDaS [ranftl2020towards], and GLIP [li2022grounded] for spatial understanding. However, existing program synthesis approaches primarily focus on single-frame reasoning. In contrast, we are interested in understanding dynamic multi-agent interactions from video sequences. We take inspiration from existing methods and apply similar program synthesis-based approaches for spatio-temporal scenario mining.
@@ 3 RefAV: Scenario Mining with Natural Language Descriptions
In this section, we present our approach for curating natural language prompts (Sec 3.1) and describe four zero-shot scenario mining baselines (Sec 3.2).
@@ 3.1 Creating RefAV
Why Use the Argoverse 2 (AV2) Sensor Dataset? The AV2 sensor dataset contains 15 – 20 second logs with synchronized sensor measurements from seven seven ring cameras and two LiDAR sensors. Moreover, the dataset includes HD maps with lane markings, crosswalk polygons, and various lane types (e.g. vehicle, bus, and bike lanes). Notably, AV2 annotates 30 object categories with track-level annotations at 10 Hz. We choose to build RefAV on top of AV2 because prior datasets like nuScenes [caesar2020nuscenes] only annotate ground-truth tracks at a lower temporal resolution (e.g. 2Hz vs. 10Hz), making it more difficult to extract fine-grained motion. We ablate this in Appendix D. In addition, KITTI [geiger2012we] and Waymo Open Dataset [sun2020scalability] only label a limited number of categories (e.g. car, pedestrian, bicycle) and do not have HD maps, making it difficult to evaluate diverse multi-agent interactions. Although AV2 annotates objects up to 150m away from the ego vehicle, we clip all object tracks at 50m. We find that current 3D perception models struggle with long-range detection and tracking [peri2023empirical], suggesting that the community is not ready to address spatio-temporal scenario mining at range.
@@
LLM-based Procedural Scenario Generation. Our key insight is that many complex scenarios (e.g. find all cars accelerating while changing lanes) can be broken down into simpler atomic actions (e.g. find cars, accelerating, and changing lanes), and a large set of such atomic actions can be composed to generate new diverse scenarios. To this end, we first define 28 atomic functions based on nuPlan’s list of planning scenarios (cf. Figure 2 left). We include a full API listing in Appendix G. To generate a new scenario definition, we provide an LLM (e.g. Claude 3.5 Sonnet) with the full API listing, along with in-context examples of real compositions. We prompt the LLM to generate permutations of the atomic functions and describe the generated code with a natural language description (cf. Figure 2 middle). We execute the generated code to filter ground-truth tracks from all logs to identify true positive matches. We aggregate all true positive log-prompt matches (>50K) and sample 8,000 true positive log-prompt pairs that maximize scenario diversity. Lastly, we randomly sample 2,000 true negative log-prompt pairs.
@@
Verifying Procedurally Generated Scenarios. Although LLM-based procedural generation allows us to generate diverse scenarios at scale, this process is not perfect and requires extensive manual validation (cf. Figure 2 right). First, we verify that the generated code aligns with the generated natural language description. We then manually review video clips of true positive log-prompt matches to ensure that the generated natural language prompt accurately describes the localized tracks. Notably, we find two common error modes. First, the LLM often reverses the relationship between objects in the code and prompt. For example, given a prompt to find bicycles in front of the car, the generated code might actually describe the scenario car in front of the bicycles. Second, we find that LLMs often generates underspecified code. For example, given the prompt bicycles traveling against traffic, the LLM correctly generates code that finds bicycles riding in the opposite direction of the ego vehicle, but it does not precisely specify that they should be on the same side of the road as the ego vehicle.
@@
Manual Scenario Annotation. Despite the versatility of our procedural generation approach, many interesting interactions cannot be easily identified by composing atomic functions. Therefore, we manually inspect all videos in the validation and test sets to identify interesting driving behaviors (cf. Fig 4). Further, we manually annotate weather (e.g. clear, cloudy, rain, snow, and fog) and lighting conditions (e.g. daylight, dusk/dawn, and night) in the validation and test sets as these attributes are relevant for establishing a safety case [lee2024typical] in all driving conditions. See Appendix B for more details about our annotation workflow.
@@
| Dataset | Base Data | Task | View | Anno. Freq. | # Expressions | # Frames | Neg. Prompts | Human Anno. | ||
| Talk2Car [deruyttere2019talk2car] | nuScenes | Detection | Front | Per Video | 11959 | 9217 | ✗ | Referring Expression | ||
| Refer-KITTI [wu2023referkitti] | KITTI | RMOT | Front | 2Hz | 818 | 6650 | ✗ | Referring Expression | ||
| nuGrounding [li2025nugrounding] | nuScenes | RMOT | 360∘ | 2Hz | 2.2M | 34149 | ✗ | Object Attribute | ||
| nuPrompt [wu2023nuprompt] | nuScenes | RMOT | 360∘ | 2Hz | 40776 | 34149 | ✗ | Object Color | ||
| nuScenes-QA [qian2024nuscenesqa] | nuScenes | VQA | 360∘ | 2Hz | 459941 | 34149 | ✓ | None | ||
| DriveLM [sima2024drivelm] | nuScenes | VQA | Front | 2Hz | 443K | 4871 | ✗ | QA Pairs | ||
| OmniDrive [wang2024omnidrive] | nuScenes | VQA | 360∘ | 2Hz | 200K | 34149 | ✗ | None | ||
| RefAV (Ours) | Argoverse 2 | Scenario Mining | 360∘ | 10Hz | 10K | 15500 | ✓ |
@@ 3.2 Scenario Mining Baselines
We present four referential tracking baselines for scenario mining, including filtering by referred class, referential track inflation, referential track classification, and referential tracking by program synthesis. We repurpose 3D trackers from winning teams in the Argoverse 2 End-to-End Forecasting Challenge.
@@
Filtering by Referred Class (RefBlind). We propose a simple “blind” baseline that does not explicitly reason about multi-agent interactions. For a given natural language prompt, we use an LLM to parse the referred object class and only keep 3D tracks from this class. For example, for the natural language query find all cars turning left, we remove all predicted tracks except cars. Interestingly, RefBlind is a surprisingly strong baseline, even compared to more complex approaches, described next.
Referential Track Inflation (RefInflate). We take inspiration from CM3D [khurana2024shelf] to create a zero-shot referential tracker using off-the-shelf 2D foundation models. First, we prompt an 2D detector with referential grounding capabilities (e.g., GroundingDINO [liu2024grounding]) with a natural language query (e.g., car turning left) to generate 2D box proposals. Next, we prompt SAM [kirillov2023segment] with the predicted 2D bounding boxes to generate high-quality instance segmentation masks. We then generate an oriented 3D cuboid using the set of LiDAR points that project to the 2D instance mask. Lastly, we associate these referred 3D detections over time using a 3D Kalman filter [weng20203d].
@@
Referential Track Classification (RefClassify). We take inspiration from ReferGPT [chamiti2025refergpt] to classify the output of off-the-shelf 3D trackers using CLIP [clip] into referred objects and other objects. For each track, we project the predicted 3D bounding box onto the 2D image plane and extract CLIP image features at each timestep. Next, we compute the cosine similarity between the per-timestep CLIP image embeddings and the CLIP text embedding of the referential prompt. We consider a track to be a referreed object if the similarity score between the image embeddings and text embedding is above a fixed threshold for more than 50% of the track length.
@@
Referential Tracking by Program Synthesis (RefProg). RefProg takes a similar approach as our LLM-based procedural scenario generator. Specifically, we prompt an LLM to generate code using an API listing of atomic actions for a given referential query. We execute the generated code on off-the-shelf 3D tracks to identify referred objects, related objects and other objects (cf. Fig 3). Despite the similarity of RefProg and our LLM-based procedural scenario generator, we highlight three key differences. First, RefProg generates code based on the referential query, while our LLM-based procedural scenario generator synthesizes the referential query based on the generated code. We posit that the role of the scenario generator is considerably easier since there are many valid natural language prompts to describe a generated program, but significantly fewer valid programs corresponding to each natural language prompt. Second, RefProg filters predicted 3D tracks, while our scenario generator uses ground-truth 3D tracks. Lastly, code from our scenario generator is verified for correctness, while code from RefProg is used without modification.

### v3

# RefAV: Towards Planning-Centric Scenario Mining

@@ Abstract
Autonomous Vehicles (AVs) collect and pseudo-label terabytes of multi-modal data localized to HD maps during normal fleet testing. However, identifying interesting and safety-critical scenarios from uncurated driving logs remains a significant challenge. Traditional scenario mining techniques are error-prone and prohibitively time-consuming, often relying on hand-crafted structured queries. In this work, we revisit spatio-temporal scenario mining through the lens of recent vision-language models (VLMs) to detect whether a described scenario occurs in a driving log and, if so, precisely localize it in both time and space. To address this problem, we introduce RefAV, a large-scale dataset of diverse natural language queries that describe complex multi-agent interactions relevant to motion planning derived from driving logs in the Argoverse 2 Sensor dataset. We evaluate several referential multi-object trackers and present an empirical analysis of our baselines. Notably, we find that naively repurposing off-the-shelf VLMs yields poor performance, suggesting that scenario mining presents unique challenges. Lastly, we discuss our recently held competition and share insights from the community. Our code and dataset are available on Github and Argoverse.
@@ 1 Introduction
Autonomous Vehicle (AV) deployment on public roads has increased significantly in recent years, with Waymo completing more than a million rides per week as of February 2025 [Korosec2025]. Despite the maturity of such autonomous ride-hailing services, AVs can still sometimes require manual interventions [korosec2025waymo, mihalascu2023cruise] and can be prone to accidents [thadani2023cruise]. Although data-driven simulators are an integral component in establishing a safety case [karnchanachari2024towards, gulino2023waymax, vasudevan2024planning, yang2023unisim], validating end-to-end autonomy with real-world operational data remains a critical part of the testing stack. However, identifying interesting and safety-critical scenarios from uncurated real-world driving logs is akin to finding a "needle in a haystack" due to the scale of data collected during normal fleet operations. In this paper, we revisit the task of spatio-temporal scenario mining with recent vision-language models (VLMs) to identify interesting multi-agent interactions using natural language (cf. Fig. 1).
Status Quo. Although language-based 3D scene understanding has been extensively studied in the context of referential multi-object tracking (RMOT) [wu2023referkitti, wu2023nuprompt, li2025nugrounding], multi-modal visual question answering (VQA) [qian2024nuscenesqa, Inoue_2024_WACV, tian2025nuscenes], and VLM-based motion planning [sima2024drivelm, mao2023gpt, shao2024lmdrive, wang2024omnidrive], we argue that spatiotemporal scenario mining presents unique challenges. RMOT extends referential grounding by associating referred objects over time. However, unlike RMOT, scenario mining does not guarantee that referred objects exist in a given driving log. Next, multi-modal VQA extends VQA with additional visual modalities like LiDAR. Although spatio-temporal scenario mining also leverages multi-modal input (e.g. RGB, LiDAR and HD maps), we require that methods output 3D tracks instead of text-based answers. Lastly, VLM-based motion planners directly estimate future ego-vehicle waypoints based on high-level language instructions and past sensor measurements. In contrast, scenario mining methods can reason over the full driving log to identify interactions between non-ego vehicles. Concretely, spatio-temporal scenario mining requires identifying if a described scenario occurs in a driving log from raw sensor measurements (e.g. RGB and LiDAR), and if so, precisely localizing all referred objects in both time and space with 3D tracks. To support this task, we propose RefAV, a large-scale dataset of diverse natural language queries designed to evaluate a model’s ability to find a visual “needle in a haystack”.
@@
Referential Tracking by Program Synthesis. Although prior referential trackers and VLMs achieve reasonable performance with simple referential prompts (e.g. find all cars), we find that such methods struggle with compositional reasoning and motion understanding (e.g. find all cars accelerating while changing lanes). To address this problem, we propose Referential Tracking by Program Synthesis (RefProg), a modular approach that combines off-the-shelf 3D tracks with LLMs. Inspired by recent work in program synthesis [surismenon2023vipergpt, gupta2023visual, marsili2025visual], our method uses an LLM to break down complex referential expressions into simpler compositional actions. Specifically, we define an API to describe hand-crafted atomic functions (e.g., turning, accelerating, changing lanes) and use an LLM to synthesize a program that composes these atomic actions corresponding to the natural language prompt. We then execute the generated program to filter off-the-shelf tracks to identify the subset of tracks that best match the described scenario (cf. Fig. 4).
@@ 2 Related Works
Visual Question Answering with VLMs has improved significantly in recent years due to large-scale multi-modal pre-training. Vision-language models (VLMs) like LLaVA [liu2023llava, liu2023improvedllava], BLIP-2 [blip2], and Qwen2.5-VL [bai2025qwen2.5vl] show strong generalization across diverse domains [balanced_vqa_v2, tian2025nuscenes, qian2024nuscenesqa]. However, state-of-the-art models still struggle with spatial reasoning and grounding [madan2024revisiting, robicheaux2025roboflow100]. To address this issue, SpatialRGPT [cheng2024spatialrgpt] and SpatialVLM [chen2024spatialvlm] generate large-scale monocular depth pseudo-labels to improve 3D reasoning. Despite the effectiveness of such methods, zero-shot prompting yields poor compositional reasoning performance. Instead, VisProg [gupta2023visual] and ViperGPT [surismenon2023vipergpt] use LLMs to generate executable code and use tools like OwlViT [minderer2022simple], CLIP [clip], MiDaS [ranftl2020towards], and GLIP [li2022grounded] for spatial understanding. However, existing program synthesis approaches primarily focus on single-frame reasoning. In contrast, we are interested in understanding dynamic multi-agent interactions from video sequences. We take inspiration from existing methods and apply similar program synthesis-based approaches for spatio-temporal scenario mining.
@@ 3 RefAV: Scenario Mining with Natural Language Descriptions
In this section, we present our approach for curating natural language prompts (Sec 3.1) and describe five zero-shot scenario mining baselines (Sec 3.2).
@@ 3.1 Creating RefAV
Why Use the Argoverse 2 (AV2) Sensor Dataset? The AV2 sensor dataset contains 15 – 20 second logs with synchronized sensor measurements from seven ring cameras and two LiDAR sensors. Moreover, the dataset includes HD maps with lane markings, crosswalk polygons, and various lane types (e.g. vehicle, bus, and bike lanes). Notably, AV2 annotates 30 object categories with track-level annotations at 10 Hz. We choose to build RefAV on top of AV2 because prior datasets like nuScenes [caesar2020nuscenes] only annotate ground-truth tracks at a lower temporal resolution (e.g. 2Hz vs. 10Hz), making it more difficult to extract fine-grained motion. We ablate this in the Appendix F. In addition, KITTI [geiger2012we] and Waymo Open Dataset [sun2020scalability] only label a limited number of categories (e.g. car, pedestrian, bicycle) and do not have HD maps, making it difficult to evaluate diverse multi-agent interactions. Although AV2 annotates objects up to 150m away from the ego vehicle, we clip all object tracks at 50m. We find that current 3D perception models struggle with long-range detection and tracking [peri2023empirical], suggesting that the community is not ready to address spatio-temporal scenario mining at range.
@@
LLM-based Procedural Scenario Generation. Our key insight is that many complex scenarios (e.g. find all cars accelerating while changing lanes) can be broken down into simpler atomic actions (e.g. find cars, accelerating, and changing lanes), and a large set of such atomic actions can be composed to generate new diverse scenarios. To this end, we first define 28 atomic functions based on nuPlan’s list of planning scenarios (cf. Figure 2 left). We include a full API listing in the Appendix J. To generate a new scenario definition, we provide an LLM (in our case, Claude 3.7 Sonnet) with the full API listing, along with in-context examples of real compositions. We prompt the LLM to generate permutations of the atomic functions and describe the generated code with a natural language description (cf. Figure 2 right). We execute the generated code to filter ground-truth tracks from all logs to identify true positive matches. We aggregate all true positive log-prompt matches (>50K) and sample 8,000 true positive log-prompt pairs that maximize scenario diversity. Lastly, we randomly sample 2,000 true negative log-prompt pairs. Our automatic scenario generation pipeline has an average success rate of 70%.
@@
Verifying Procedurally Generated Scenarios. Although LLM-based procedural generation allows us to generate diverse scenarios at scale, this process is not perfect and requires extensive manual validation (cf. Figure 2 right). First, we verify that the generated descriptions match the generated code. We then manually review video clips of true positive log-prompt matches to ensure that the generated natural language description accurately describes the localized tracks. Notably, we find two common error modes. First, the LLM often defines the prototypical case of a scenario, but misses or incorrectly includes edge cases due to under or overspecification. For example, given a description car following a bicyclist, the generated code identifies instances where a moving car is behind a moving bike while traveling in the same road lane. This definition includes the false positive edge case where the car and bicyclist are traveling in opposite directions. Second, we find that LLMs often reverse the relationship of referred and related objects. For example, given the description bicycles in front of the car, the LLM generates code that corresponds to car behind the bicycles. Based on the videos, human annotators make the necessary changes to the scenario definition, and the modified code is re-executed across all logs. It takes about 3 minutes to verify and edit each scenario definition. In total, we spent 200 hours verifying scenario definitions.
@@
Manual Scenario Annotation. Despite the versatility of our procedural generation approach, many interesting interactions cannot be easily identified by composing atomic functions. Therefore, we manually inspect all videos in the validation and test sets to identify interesting driving behaviors (cf. Fig 3). Further, we manually annotate weather (e.g. clear, cloudy, rain, snow, and fog) and lighting conditions (e.g. daylight, dusk/dawn, and night) in the validation and test sets as these attributes are relevant for establishing a safety case [lee2024typical] in all driving conditions. See the Appendix D for more details about our annotation workflow.
@@
| Dataset | Base Data | Task | View | Anno. Freq. | # Expressions | # Frames | Neg. Prompts | Human Anno. | ||||
| Talk2Car [deruyttere2019talk2car] | nuScenes | Detection | Front | Per Video | 12k | 9k | ✗ | Referring Expression | ||||
| nuScenes-QA [qian2024nuscenesqa] | nuScenes | VQA | 360∘ | 2Hz | 460k | 34k | ✓ | None | ||||
| DriveLM [sima2024drivelm] | nuScenes | VQA | 360∘ | 2Hz | 443k | 5k | ✗ | QA Pairs | ||||
| OmniDrive [wang2024omnidrive] | nuScenes | VQA | 360∘ | 2Hz | 200k | 34k | ✗ | None | ||||
| Refer-KITTI [wu2023referkitti] | KITTI | RMOT | Front | 2Hz | 818 | 7k | ✗ | Referring Expression | ||||
| nuGrounding [li2025nugrounding] | nuScenes | RMOT | 360∘ | 2Hz | 2.2M | 34k | ✗ | Object Attribute | ||||
| nuPrompt [wu2023nuprompt] | nuScenes | RMOT | 360∘ | 2Hz | 407k | 34k | ✗ | Object Color | ||||
| RefAV | Argoverse 2 |
360∘ | 10Hz | 10k | 155k | ✓ |
@@ 3.2 Scenario Mining Baselines
We present five referential tracking baselines for scenario mining, including filtering by referred class, ReferGPT [chamiti2025refergpt], image embedding similarity, LLM APIs as a black box, and referential tracking by program synthesis. We repurpose 3D trackers from the AV2 End-to-End Forecasting Challenge
@@
Filtering by Referred Class. We propose a simple “blind” baseline that does not explicitly reason about multi-agent interactions. For a given natural language prompt, we use an LLM to parse the referred object class and only keep 3D tracks from this class. For example, for the natural language query find all cars turning left, we remove all predicted tracks except cars.
ReferGPT. ReferGPT [chamiti2025refergpt] classifies the output of off-the-shelf 3D trackers into referred objects and other objects using CLIP text embeddings. First, ReferGPT projects the predicted 3D bounding box of each track onto the 2D image plane at each timestamp to get an image crop of each object. Next, ReferGPT parses the object’s coordinates, velocity, yaw, yaw rate, and distance from the ego-vehicle. Using both the object image crop and bounding box descriptor, ReferGPT uses a VLM to generate a descriptive caption. Finally, ReferGPT uses the text cosine-similarity between the generated caption and referential prompt to score the relevance of the object track. We use GPT-5-mini as our captioning VLM and SigLIP2 [tschannen2025siglip2multilingualvisionlanguage] to compute text embeddings. We make several modifications to ReferGPT to adapt it for the spatio-temporal scenario mining problem. Since an object may be visible in multiple cameras at a single timestamp, we only caption the 2D bounding box with the largest area at each timestamp. ReferGPT selects referred objects by clustering tracks with the highest similarity scores. Since RefAV contains also many negative prompts, we consider a track to be a referred object if the similarity score between the caption text embeddings and prompt text embedding are above a fixed threshold for more than 50% of the predicted track length.
@@
Image-Embedding Similarity. We take inspiration from ReferGPT to classify the output of off-the-shelf 3D trackers into referred objects and other objects by directly using CLIP [clip] image features. We extract the track image crops in the same way as ReferGPT and compute CLIP image features for each track at each timestep. Next, we compute the cosine similarity between the per-timestep CLIP image embeddings and the CLIP text embedding of the referential prompt. We consider a track to be a referred object if the similarity score between the image and text embeddings are above a fixed threshold for more than 50% of the track length.
LLM API as a Black Box. We evaluate GPT-5 as a black-box scenario mining algorithm. For each log in AV2, we ask the OpenAI Responses API to identify which predictions match the referential prompt, and provide map data, city to ego-vehicle transforms, and 3D track information from an off-the-shelf tracker. GPT-5 responds with a CSV listing all referred objects from the initial set of predictions. Interestingly, analyzing the thinking traces of GPT-5 shows that it writes code from scratch to parse the input files and reason about multi-agent interactions.
@@
Referential Tracking by Program Synthesis (RefProg). RefProg takes a similar approach as LLM API as a Black Box, but provides more code scaffolding to improve accuracy. Specifically, we prompt an LLM to generate code using an API listing of atomic actions for a given referential query. We execute the generated code on off-the-shelf 3D tracks to identify referred objects, related objects and other objects (cf. Fig 4). In addition to using API listings for estimating an object’s state and relationships from 3D tracks, RefProg is given access to visual tools such as SigLIPv2 to identify visual attributes of tracked objects. Despite the similarity of RefProg and our LLM-based procedural scenario generator, we highlight two key differences. First, RefProg generates code based on the referential query, while our LLM-based procedural scenario generator synthesizes the referential query based on the generated code. We posit that the role of the scenario generator is considerably easier since there are many valid natural language prompts to describe a generated program, but significantly fewer valid programs corresponding to each natural language prompt. Lastly, code from our scenario generator is extensively modified by human annotators for correctness, while code from RefProg is executed without verification. Notably, we also evaluate RefProg on nuPrompt to demonstrate that the atomic functions generalize beyond RefAV.